# Adaptation of Agentic AI

### v2

@@ 7 Applications
The rapid advancement of agentic AI systems has led to their adoption across a growing range of scientific and engineering domains. In this section, we organize representative applications by discipline. Specifically, we categorize these applications into the following major areas: General Science, such as Deep Research7.1); Computer Science, where agents augment or automate processes in Software Development7.2) and Computer Use7.3); and Biomedicine, where agents accelerate research in drug discovery and development7.4).
@@ 7.1 Deep Research
Deep research systems represent the emerging class of AI-powered applications designed to automate end-to-end scientific investigation by integrating large language models (LLMs), advanced retrieval, and autonomous reasoning [6]. OpenAI’s DeepResearch [194] is a prominent example, featuring a multi-step reasoning workflow that conducts iterative search, validation, and synthesis. Similar paradigms have been adopted in recently announced systems such as Claude’s deep-search capabilities [195] and Google’s Gemini-based research agents [196]. The defining distinction of deep research systems, compared to general-purpose AI agents, is their dual adaptation in both agent reasoning and scientific tool integration.
Agent adaptation
Deep research systems require sophisticated agentic workflows capable of decomposing complex scientific questions into structured research plans. This includes: (1) adapting underlying LLMs toward long-context reasoning, hypothesis refinement, and multi-step self-critique, (2) orchestrating multiple agents to collaborate hierarchically, e.g., for literature review, data interpretation, and conclusion synthesis, and (3) maintaining persistent memory and knowledge tracking across long investigative trajectories. These adaptations enable agents not only to respond to queries but to behave as autonomous researchers navigating the breadth of scientific knowledge.
Tool adaptation
Reliable research requires grounded evidence. To address hallucination and improve informativeness, deep research agents must incorporate diverse tools that provide direct access to external knowledge, including: (1) structured retrieval interfaces to literature databases (e.g., PubMed, arXiv), (2) web navigation tools for interacting with scientific resources, (3) and modular computational utilities for data analysis and visualization. Recent advances further enhance tool adaptation through learning-based retrieval modules, such as DeepRetrieval [21] and s3 [27], which boost accuracy in real-time information gathering, especially when operating atop proprietary models that cannot be fine-tuned.
Toward domain-specialized deep research
While current systems are primarily built on generic corpora and may struggle with nuanced expert-level inquiries, the paradigm naturally extends to specialized scientific fields. Future development will increasingly involve integrating: domain knowledge bases and ontologies, validated bioinformatics and biomedical computation tools, field-specific safety, reliability, and evaluation protocols.
Such advancements are expected to transform deep research systems from broad “knowledge navigators” into expert collaborators for vertical domains like medicine, materials science, and drug development, guiding researchers from problem conception to actionable discoveries.
@@ 7.2 Software Development
AI-assisted software development represents one of the most technically demanding and economically significant domains for agentic AI systems. Unlike conventional code completion systems, software development agents are designed to autonomously navigate multi-stage engineering workflows, including requirement interpretation, code generation, debugging, testing, and deployment, within real development environments. Modern systems, including Cursor [197], Claude Code [198], and OpenAI’s CodeX [199], exemplify this shift from passive code assistants toward interactive, full-cycle programming agents capable of understanding project context and performing tool-mediated reasoning. To evaluate these capabilities, the SWE-Bench benchmark [200] has emerged as a representative testing suite that measures an agent’s ability to autonomously fix real-world software bugs in open-source repositories by reading, editing, and validating code through continuous integration workflows.
@@
Building effective software agents in this setting requires both agent adaptation, which strengthens reasoning, planning, and self-verification across complex development pipelines, and tool adaptation, which integrates and evolves the surrounding development ecosystem such as compilers, debuggers, and test frameworks.
Agent adaptation
Agent adaptation in software development focuses on enhancing model reasoning and autonomy across complex multi-stage engineering workflows. Recent systems train agents directly from interaction trajectories within real or simulated development environments.
Tool adaptation
Tool adaptation in this domain involves evolving the software ecosystem itself to improve the reliability, responsiveness, and contextual integration of tools that agents depend on for code execution, testing, and evaluation. Instead of merely wrapping existing IDE functionalities, modern systems are increasingly training tools for agents to optimize for usability and feedback efficiency. A representative example is Cursor’s Tab-RL framework [203], which applies reinforcement learning to refine the editor’s tab completion behavior based on real-world user interactions, aligning the tool’s interface dynamics with agent and developer preferences. A more advanced example is SWE-Grep [193], a specialized sub-agent trained using reinforcement learning for fast, multi-turn, and highly parallel context retrieval. By delegating code search to this T2-style tool, the main agent’s context window is conserved and protected from irrelevant “context pollution”, allowing it to focus on higher-level reasoning. More broadly, this category of tool adaptation includes the automated creation or refinement of compilers, debuggers, and linters that provide structured feedback loops for agents.
@@ 7.3 Computer Use
Computer-use agents represent an emerging class of multimodal AI systems capable of autonomously operating computers and software environments through direct interaction with graphical user interfaces (GUIs). Rather than relying on predefined APIs or task-specific integrations, these agents perceive screens as visual input, reason about interface elements such as buttons, menus, and text fields, and execute actions using a virtual keyboard and mouse—closely mirroring human computer operation. A recent example is OpenAI’s Computer-Using Agent (CUA) [204], which combines vision-based perception with reinforcement learning to navigate complex digital environments. This paradigm signifies a step toward generalized digital intelligence, where agents can perform diverse tasks, such as information retrieval, document editing, and software automation, directly within existing human-designed computing ecosystems.
@@
Representative benchmarks for this paradigm include OSWorld [205], WebArena [206], VisualWebArena [207], AppWorld [208], WebVoyager [209], and $\tau$-bench [210] which evaluate an agent’s ability to perceive, reason, and act across diverse digital environments ranging from full operating systems to real-world web interfaces. Achieving reliable and efficient performance in computer-use scenarios requires adaptation from the agent and tool levels.
Agent adaptation
@@
To handle complex computer-use tasks, agent adaptation plays a crucial role in equipping models with new knowledge and operational skills beyond those learned from general-purpose pretraining. Such adaptation often involves exposing the agent to realistic or synthesized trajectories of GUI-based interactions, enabling it to acquire procedural competence in perceiving interface states, reasoning over visual elements, and executing multi-step actions. A representative example is OpenCUA [211], which shows how large-scale, GUI-centric data can significantly improve an agent’s computer-use abilities. By collecting human demonstrations across diverse operating systems and applications, and converting them into state–action trajectories with reflective reasoning, OpenCUA provides agents with realistic exposure to interface dynamics. Another approach to agent adaptation is explored in AgentTrek [212], which takes a different approach to agent adaptation by synthesizing training trajectories from web tutorials instead of relying on human demonstrations. It converts tutorial text into step-by-step goals and has a VLM agent execute them in real environments, keeping only correct trajectories through automatic evaluation. This scalable data generation helps agents acquire new skills and interface patterns at low cost, showing that synthesized trajectories can effectively support GUI-agent adaptation.
Tool adaptation
@@
Complementary to agent-level adaptation, tool adaptation aims to enhance the tools and interfaces that agents rely on, enabling them to become more adaptive, context-aware, and synergistic with the agent’s reasoning process. Instead of modifying model parameters, these approaches update or expand the tool’s experience pool, memory, or contextual representations to better support long-horizon interaction and dynamic task requirements. A representative example is Agentic Context Engineering (ACE) [213], which treats evolving contexts as structured playbooks that accumulate, refine, and organize strategies for tool use. By continuously curating and updating contextual knowledge through execution feedback, ACE effectively adapts the operational layer of tools—reducing rollout latency and improving alignment with the agent’s decision-making. Such approaches highlight a broader trend: as agents become more capable, the tools they employ must likewise evolve, incorporating persistent memory and adaptive control mechanisms to ensure seamless collaboration in open computer-use environments.
@@ 7.4 Drug Discovery and Development
LLM-empowered AI agents are rapidly transforming biomedical research and, by extension, the entire drug discovery and development pipeline [5]. Modern systems increasingly integrate both agent adaptation (e.g., fine-tuning LLMs and designing agentic workflows) and tool adaptation (e.g., incorporating domain-specific databases, scientific software, and retrieval components) [8]. These two forms of adaptation are fundamentally complementary: agent adaptation improves reasoning and procedural reliability, whereas tool adaptation equips agents with practical scientific capabilities. Below, we describe representative advances that emphasize one aspect more than the other, while acknowledging that most systems combine both.
Agent adaptation for drug discovery
@@
Agent adaptation for drug development
Clinical research is a central component of drug development, and agents are increasingly tailored to literature analysis, patient recruitment, and trial design. TrialMind adapts LLM agents for biomedical evidence retrieval by integrating medical guidelines and structured access to clinical-trial and publication databases to support search, screening, and data extraction [7]. LEADS builds on this by further training the model with curated literature corpora to improve agent-driven evidence discovery [217]. Similarly, TrialGPT operationalizes guideline-based patient-to-trial matching through a reason-and-retrieve workflow [12]. For upstream design tasks, TrialGenie leverages multi-agent collaboration to parse historical trial documents and generate analytical code for real-world datasets [218].
Tool adaptation
Tool adaptation has progressed equally rapidly, driven by the need to support diverse scientific tasks. Tools such as SyntheMol and related frameworks integrate ML-based molecular property predictors as reward functions to steer generative models toward biologically desirable compounds [219, 220]. ToolUniverse focuses on scalable scientific tool creation: a discovery module constructs tools from natural language specifications, and an optimizer iteratively refines them before incorporation into a shared library for custom agent assembly [221]. Biomni takes a complementary approach by manually mining biomedical literature to curate a high-quality tool repository, which can be dynamically injected into agent workflows [222]. Meanwhile, STELLA proposes a self-evolving paradigm where an expanding Template Library captures reasoning strategies and a Tool Ocean continuously grows as a tool-creation agent autonomously discovers and integrates new bioinformatics utilities [223].
@@ 8 Opportunities
This paper has systematically categorized the landscape of agentic AI adaptation into four key paradigms: (A1) Agent Adaptation with Tool Execution Signal, (A2) Agent Adaptation with Agent Output Signal, (T1) Agent-Agnostic Tool Adaptation, and (T2) Agent-Supervised Tool Adaptation. These paradigms provide a crucial framework for organizing and understanding current methods. However, their true value lies in illuminating the path forward. The separation of agent and tool adaptation, while analytically useful, is largely a construct of the field’s nascent stage; the future of capable, robust, and efficient agentic AI will almost certainly be defined by their synthesis.
This section functions as a forward-looking roadmap, identifying some critical and interdependent opportunities for future research that emerge directly from our taxonomy. These opportunities represent the next frontier of agentic AI, moving from static, monolithic adaptation toward dynamic, co-adaptive, and federated systems.
@@ 8.1 Co-Adaptation
The taxonomy presented in this paper (A1/A2 vs. T1/T2) is a necessary simplification, organizing the field by its dominant locus of optimization: either the agent or its tools. The most significant and challenging opportunity for the next decade of research is to dissolve this boundary and develop unified agent-tool co-adaptation frameworks.
@@
Such a framework implies a complex, bi-level optimization problem, formally $max_{\mathcal{A},\mathcal{T}}\mathcal{O}(\mathcal{A},\mathcal{T})$, where the agent’s policy ($\mathcal{A}$) and the tool’s internal parameters ($\mathcal{T}$) are adapted simultaneously within the same learning loop. This represents a fundamental departure from current paradigms, which almost universally rely on freezing one component to provide a stable learning target for the other (e.g., $\mathcal{A}_{\text{frozen}}$ in T2, or $\mathcal{T}_{\text{frozen}}$ in A1/A2).
This is not an entirely new problem, and researchers can draw deep conceptual inspiration from several established fields:
@@
  • Co-evolutionary Algorithms. Classic work in evolutionary computation has long studied how two or more interacting populations—such as hosts vs. parasites or predators vs. prey—apply reciprocal selection pressures that drive arms races, emergent structure, and increasingly sophisticated strategies. Hillis [224] introduced the seminal host–parasite model, showing that co-evolving adversarial test cases can significantly improve solution robustness. Subsequent work on competitive co-evolution explored dynamics such as disengagement, cycling, and evolutionary complexification [225]. Other lines of research developed cooperative multi-population architectures in which sub-components co-adapt to form joint solutions [226]. Comprehensive surveys [227] situate these approaches within a broader taxonomy of competitive and cooperative CEAs. In our setting, we can view the agent $\mathcal{A}$ and its tool $\mathcal{T}$ as two interdependent populations evolving on a shared fitness landscape, allowing reciprocal adaptation, arms-race dynamics, and emergent specialization to arise naturally from their interaction.
@@
  • Multi-Agent Systems. A complementary line of work emerges from multi-agent reinforcement learning, where each agent learns in a non-stationary environment induced by other concurrently learning agents. Foundational surveys [228, 229, 230] describe how decentralized learners must cope with shifting policies, partial observability, and strategic coupling, challenges that closely mirror those of agent–tool co-adaptation. Classic problems such as equilibrium selection, credit assignment, and coordination under changing partner behaviors have led to techniques including opponent modeling, joint-policy search, centralized training with decentralized execution (CTDE), and communication-based coordination. In our context, viewing $\mathcal{A}$ and $\mathcal{T}$ as a two-agent partially cooperative system highlights the need for algorithms that stabilize learning under mutual adaptation, prevent non-stationarity-induced divergence, and support the emergence of complementary capabilities rather than competitive oscillations.
A primary technical barrier to effective co-adaptation is the intractable credit assignment problem. When an agentic system fails at a complex task, the source of the failure is inherently ambiguous. Consider a system in which an A2-style planner invokes a T2-style search subagent (e.g., an s3-like searcher). If the final answer is incorrect, which component is responsible?
Nascent research is beginning to address fragments of this joint-optimization challenge. MATPO (Multi-Agent Tool-Integrated Policy Optimization) [231] proposes a “principled credit assignment mechanism” for jointly training planner and worker agents. However, in its current form, these “agents” correspond to distinct prompt roles instantiated within a single LLM, rather than heterogeneous models. Other work studies joint refinement of agent prompts and tool specifications [232]. The true opportunity lies in extending these ideas to distributed, heterogeneous systems in which the agent $\mathcal{A}$ and tools $\mathcal{T}$ are distinct learning entities. This may require importing architectures from multi-agent RL (such as centralized-critic–decentralized-actor methods [233]) to enable principled credit allocation over an interconnected agent–tool graph.
@@
A deeper difficulty involves the Stability–Plasticity Dilemma. Co-adaptation aims for $\mathcal{A}$ and $\mathcal{T}$ to become mutually optimized, yet in a joint-learning framework, $\mathcal{A}$ is adapting to a $\mathcal{T}$ that is itself changing. As established in the study of complex adaptive systems, such non-stationarity can induce chaotic or unstable dynamics [234]. The system may enter a “Red Queen” regime in which $\mathcal{A}$ and $\mathcal{T}$ continually adjust to each other’s most recent changes without increasing overall performance, or may even collapse into degenerate policies. Conversely, premature convergence may cause the system to “lock in” a brittle, suboptimal agent–tool interface, losing the plasticity required for generalization.
A key research direction is the development of pacemaker mechanisms that regulate the relative learning rates of agents and tools, or the use of evolutionary game-theoretic analyses to guarantee convergence toward stable symbiotic equilibria [234]. These mechanisms will be essential for enabling reliable, scalable co-adaptation in next-generation agentic AI systems.

### v3

@@ 8 Applications
Agentic AI systems have been adopted across a growing range of scientific and engineering domains. The following subsections organize representative applications by discipline and connect each to the adaptation paradigms (A1/A2/T1/T2) developed in this paper. We categorize these applications into the following areas: General Science, such as Deep Research8.1); Computer Science, where agents augment or automate processes in Software Development8.2) and Computer Use8.3); and Biomedicine, where agents accelerate research in drug discovery and development8.4).
Across these domains, the dominant adaptation paradigm varies with the availability of verifiable feedback and the cost of agent retraining. Table 9 profiles each domain by its dominant paradigm, key bottleneck, and a representative system.
@@ 8.1 Deep Research
Deep research systems automate end-to-end scientific investigation by integrating large language models (LLMs), advanced retrieval, and autonomous reasoning [6]. OpenAI’s DeepResearch [302] is a prominent example, featuring a multi-step reasoning workflow that conducts iterative search, validation, and synthesis. Similar paradigms have been adopted in recently announced systems such as Claude’s deep-search capabilities [303] and Google’s Gemini-based research agents [304]. Their defining characteristic, compared to general-purpose AI agents, is dual adaptation in both agent reasoning and scientific tool integration.
Agent adaptation.
Deep research systems require agentic workflows that decompose complex scientific questions into structured research plans: (1) adapting LLMs toward long-context reasoning, hypothesis refinement, and multi-step self-critique, (2) orchestrating multiple agents to collaborate hierarchically for literature review, data interpretation, and conclusion synthesis, and (3) maintaining persistent memory and knowledge tracking across long investigative trajectories.
Tool adaptation.
To address hallucination and improve informativeness, deep research agents incorporate diverse tools that provide direct access to external knowledge: (1) structured retrieval interfaces to literature databases (e.g., PubMed, arXiv), (2) web navigation tools for interacting with scientific resources, and (3) modular computational utilities for data analysis and visualization. Learning-based retrieval modules such as DeepRetrieval [22] and s3 [28] further advance tool adaptation, boosting accuracy in real-time information gathering, especially atop proprietary models that cannot be fine-tuned.
Toward domain-specialized deep research.
Current systems are primarily built on generic corpora and may struggle with nuanced expert-level inquiries. Extension to specialized scientific fields will involve integrating domain knowledge bases and ontologies, validated bioinformatics and biomedical computation tools, and field-specific safety, reliability, and evaluation protocols.
Such integration would enable deep research systems to function as expert collaborators for vertical domains such as medicine, materials science, and drug development.
Paradigm mapping.
Deep research exemplifies the complementary use of multiple adaptation paradigms. The core reasoning agent is typically adapted via A2 (optimizing end-to-end research quality), while retrieval and search tools are adapted via T2 (training search subagents like s3 [28] under frozen-agent supervision) or deployed as T1 components (pre-trained dense retrievers). The absence of deterministic execution signals (unlike code compilation) makes pure A1 adaptation less natural in this domain, though retrieval-specific metrics (Recall@$K$) can provide A1-style feedback for the search component.
@@ 8.2 Software Development
AI-assisted software development is a major application area for agentic AI. Unlike conventional code completion systems, software development agents autonomously navigate multi-stage engineering workflows (requirement interpretation, code generation, debugging, testing, and deployment) within real development environments. Modern AI-assisted development tools and agents such as Cursor [305], Claude Code [306], and OpenAI’s CodeX [307] exemplify this shift from passive code completion toward interactive, full-cycle programming capable of understanding project context and performing tool-mediated reasoning. To evaluate these capabilities, the SWE-Bench benchmark [208] serves as a representative testing suite that measures an agent’s ability to autonomously fix real-world software bugs in open-source repositories by reading, editing, and validating code through continuous integration workflows.
@@
Building effective software agents requires both agent adaptation (strengthening reasoning, planning, and self-verification across complex development pipelines) and tool adaptation (integrating and evolving the surrounding development ecosystem such as compilers, debuggers, and test frameworks).
Agent adaptation.
Agent adaptation in software development improves model reasoning and autonomy across multi-stage engineering workflows. The availability of deterministic execution signals (test-suite pass rates, compilation success, CI/CD pipeline outcomes) makes software development well suited for A1-style adaptation, where agents are optimized on tool-execution feedback [21, 66, 67]. A2-style adaptation applies when agents are evaluated on holistic task completion (e.g., resolving a full GitHub issue), which requires reasoning about when to read code, what to edit, and how to validate changes.
Tool adaptation.
Tool adaptation in this domain involves evolving the software ecosystem to improve the reliability, responsiveness, and contextual integration of tools that agents depend on for code execution, testing, and evaluation. A representative example is Cursor’s Tab-RL framework [310], which applies reinforcement learning to refine the editor’s tab completion behavior based on real-world user interactions, aligning the tool’s interface dynamics with agent and developer preferences. A more advanced example is SWE-Grep [243], a specialized subagent trained using reinforcement learning for fast, multi-turn, and highly parallel context retrieval. Delegating code search to this T2-style tool conserves the main agent’s context window and protects it from irrelevant “context pollution.” This category of tool adaptation also includes the automated creation or refinement of compilers, debuggers, and linters that provide structured feedback loops for agents.
@@ 8.3 Computer Use
Computer-use agents are multimodal AI systems that autonomously operate computers through direct interaction with graphical user interfaces (GUIs). These agents perceive screens as visual input, reason about interface elements (buttons, menus, text fields), and execute actions using a virtual keyboard and mouse. OpenAI’s Computer-Using Agent (CUA) [311], which combines vision-based perception with reinforcement learning, exemplifies this paradigm.
@@
Representative benchmarks for this paradigm include OSWorld [256], WebArena [207], VisualWebArena [291], AppWorld [312], WebVoyager [313], and $\tau$-bench [257] which evaluate an agent’s ability to perceive, reason, and act across diverse digital environments ranging from full operating systems to real-world web interfaces. Reliable performance in computer-use scenarios requires adaptation at both the agent and tool levels.
Agent adaptation.
@@
Agent adaptation equips models with operational skills beyond those learned from general-purpose pre-training by exposing them to realistic or synthesized trajectories of GUI-based interactions. A representative example is OpenCUA [314], which shows how large-scale, GUI-centric data can improve an agent’s computer-use abilities. By collecting human demonstrations across diverse operating systems and applications, and converting them into state-action trajectories with reflective reasoning, OpenCUA provides agents with realistic exposure to interface dynamics. AgentTrek [315] takes a complementary approach by synthesizing training trajectories from web tutorials instead of relying on human demonstrations. It converts tutorial text into step-by-step goals and has a VLM agent execute them in real environments, keeping only correct trajectories through automatic evaluation. The resulting data generation scales at low cost, showing that synthesized trajectories can effectively support GUI-agent adaptation.
Tool adaptation.
@@
Tool adaptation improves the tools and interfaces that agents rely on. Instead of modifying model parameters, these approaches update or expand the tool’s experience pool, memory, or contextual representations to better support long-horizon interaction. CUA-Skill [316] provides a concrete skill-centric example: it builds a reusable computer-use skill base with parameterized execution and composition graphs, allowing an agent to retrieve skills, instantiate arguments, and recover from failures across Windows applications. The design shows that, in GUI environments, a large part of the adaptation burden can sit in the skill substrate rather than in the agent weights alone. A representative example of adaptive context management is Agentic Context Engineering (ACE) [317], which treats evolving contexts as structured playbooks that accumulate, refine, and organize strategies for tool use. By continuously curating and updating contextual knowledge through execution feedback, ACE adapts the operational layer of tools, reducing rollout latency and improving alignment with the agent’s decision-making.
Paradigm mapping.
Computer use requires multimodal adaptation: the agent must ground its actions in visual perception (screenshots, UI layouts) rather than purely textual signals. A1-style adaptation is less natural here, as verifiable execution signals are harder to define (there is no “test suite” for clicking the right button). The dominant paradigm is A2-style adaptation with holistic task-completion rewards, supplemented by T2-style tool adaptation for persistent memory and context management. T1 tools (pre-trained vision models like CLIP [131] and SAM [132]) serve as plug-and-play perception modules.
More broadly, as agents become more capable, the tools they employ must likewise evolve, incorporating persistent memory and adaptive control mechanisms to support effective collaboration in open computer-use environments.
@@ 8.4 Drug Discovery and Development
LLM-based AI agents are increasingly applied across the drug discovery pipeline [5]. Modern systems integrate both agent adaptation (fine-tuning LLMs and designing agentic workflows) and tool adaptation (incorporating domain-specific databases, scientific software, and retrieval components) [8]. Agent adaptation improves reasoning and procedural reliability, whereas tool adaptation equips agents with practical scientific capabilities.
Agent adaptation for drug discovery.
@@
Agent adaptation for drug development.
Clinical research involves literature analysis, patient recruitment, and trial design—tasks that differ in automation readiness. Evidence retrieval (TrialMind [7], LEADS [321]) admits structured supervision via citation recall, making it amenable to A1-style feedback. Patient-to-trial matching (TrialGPT [12]) is inherently holistic (guideline-based eligibility criteria resist decomposition into verifiable sub-signals), favoring A2. Upstream trial design (TrialGenie [322]), which uses multi-agent collaboration to parse historical protocols and generate analytical code, combines both: code-execution signals (A1) and overall protocol quality (A2).
Tool adaptation.
[…]
Paradigm mapping.
Drug discovery and development benefits from all four adaptation paradigms. A1 applies when agents interact with computational chemistry tools that provide verifiable outputs (e.g., docking scores, molecular property predictions). A2 governs the higher-level research workflow (e.g., hypothesis generation, literature synthesis) where holistic quality matters. T1 tools are abundant: AlphaFold2 [139], ESMFold [140], and molecular property predictors are pre-trained independently and used as plug-and-play components. T2 adaptation is exemplified by systems like STELLA and ToolUniverse, where the tool ecosystem evolves under agent supervision. The long feedback loops inherent in wet-lab validation (weeks to months) make this domain particularly challenging, as the reward signal is sparse and delayed compared to code execution or retrieval.
Table 9: Adaptation profile of each application domain: which paradigm dominates, what bottleneck limits further progress, and a representative system illustrating the dominant paradigm.
Domain Dominant Paradigm Key Bottleneck Representative System
Deep Research T2 (search/planning subagent) No deterministic execution signal; holistic quality is hard to verify AgentFlow [52]
Software Dev. A1 (execution RL) Long-horizon credit assignment across multi-file edits RLEF [21]
Computer Use A2 (task completion) Visual grounding; no “test suite” for GUI actions OpenCUA [314]
Drug Discovery T1 (plug-in models) Sparse, delayed wet-lab feedback (weeks–months) AlphaFold2 [139]
@@ 9 Opportunities
The preceding sections organized agentic AI adaptation into four paradigms: (A1) Agent Adaptation with Tool Execution Signal, (A2) Agent Adaptation with Agent Output Signal, (T1) Agent-Agnostic Tool Adaptation, and (T2) Agent-Supervised Tool Adaptation. These paradigms provide a framework for organizing current methods, but their value also lies in clarifying the path forward. The separation of agent and tool adaptation, while analytically useful, reflects the field’s present state of development; capable, robust, and efficient agentic AI will likely require their synthesis.
We identify opportunities for future research that emerge from our taxonomy, organized from single-component optimization toward joint agent-tool learning, continual adaptation, safety, and efficiency.
@@ 9.1 Co-Adaptation
The taxonomy presented in this paper (A1/A2 vs. T1/T2) is a necessary simplification, organizing the field by its dominant locus of optimization: either the agent or its tools. The central challenge for the next stage of research is to dissolve this boundary and develop unified agent-tool co-adaptation frameworks.
@@
Such a framework implies a complex, bi-level optimization problem, where the agent’s policy ($\mathcal{A}$) and the tool’s parameters ($\mathcal{T}$) are adapted simultaneously. Naïvely stated as $\max_{\mathcal{A},\mathcal{T}}\mathcal{O}(\mathcal{A},\mathcal{T})$, the problem has a bi-level structure because $\mathcal{A}$’s optimal policy depends on $\mathcal{T}$ and vice versa. The formulation departs from current paradigms, which almost universally rely on freezing one component to provide a stable learning target for the other (e.g., $\mathcal{A}_{\text{frozen}}$ in T2, or $\mathcal{T}_{\text{frozen}}$ in A1/A2).
Related problems have been studied in several established fields:
@@
  • Co-evolutionary Algorithms. Classic work in evolutionary computation has studied how two or more interacting populations (such as hosts vs. parasites or predators vs. prey) apply reciprocal selection pressures that drive arms races, emergent structure, and progressively more complex strategies. Hillis [328] introduced the seminal host-parasite model, showing that co-evolving adversarial test cases can improve solution robustness. Subsequent work on competitive co-evolution explored dynamics such as disengagement, cycling, and evolutionary complexification [329]. Other lines of research developed cooperative multi-population architectures in which subcomponents co-adapt to form joint solutions [330]. Comprehensive surveys [331] situate these approaches within a broader taxonomy of competitive and cooperative CEAs. In our setting, we can view the agent $\mathcal{A}$ and its tool $\mathcal{T}$ as two interdependent populations evolving on a shared fitness landscape, where reciprocal selection pressures can drive emergent specialization.
@@
  • Multi-Agent Systems. A complementary line of work emerges from multi-agent reinforcement learning, where each agent learns in a non-stationary environment induced by other concurrently learning agents. Foundational surveys [332, 293, 333] describe how decentralized learners must cope with shifting policies, partial observability, and strategic coupling, challenges that closely mirror those of agent-tool co-adaptation. Classic problems such as equilibrium selection, credit assignment, and coordination under changing partner behaviors have led to techniques including opponent modeling, joint-policy search, centralized training with decentralized execution (CTDE), and communication-based coordination. In our context, viewing $\mathcal{A}$ and $\mathcal{T}$ as a two-agent partially cooperative system highlights the need for algorithms that stabilize learning under mutual adaptation, prevent non-stationarity-induced divergence, and support the emergence of complementary capabilities rather than competitive oscillations.
@@
The first technical barrier to effective co-adaptation is the intractable credit assignment problem. When an agentic system fails at a complex task, the source of the failure is ambiguous. Consider a system in which an A2-style planner invokes a T2-style search subagent (e.g., an s3-like searcher). If the final answer is incorrect, which component is responsible?
Recent work has begun to address fragments of this joint-optimization challenge. MATPO (Multi-Agent Tool-Integrated Policy Optimization) [334] proposes a credit assignment mechanism for jointly training planner and worker agents. However, in its current form, these “agents” correspond to distinct prompt roles instantiated within a single LLM, rather than heterogeneous models. Other work studies joint refinement of agent prompts and tool specifications [335]. The open challenge is to extend these ideas to distributed, heterogeneous systems in which the agent $\mathcal{A}$ and tools $\mathcal{T}$ are distinct learning entities. Addressing the challenge may require importing architectures from multi-agent RL (such as centralized-critic, decentralized-actor methods [336]) to enable principled credit allocation over an interconnected agent-tool graph.
A second difficulty involves the stability-plasticity dilemma. Co-adaptation aims for $\mathcal{A}$ and $\mathcal{T}$ to become mutually optimized, yet in a joint-learning framework, $\mathcal{A}$ is adapting to a $\mathcal{T}$ that is itself changing. As established in the study of complex adaptive systems, such non-stationarity can induce chaotic or unstable dynamics [337]. The system may enter a “Red Queen” regime in which $\mathcal{A}$ and $\mathcal{T}$ continually adjust to each other’s most recent changes without increasing overall performance, or may even collapse into degenerate policies. Conversely, premature convergence may cause the system to lock in a brittle, suboptimal agent-tool interface, losing the plasticity required for generalization.
One concrete direction is the development of pacemaker mechanisms that regulate the relative learning rates of agents and tools, or the use of evolutionary game-theoretic analyses to guarantee convergence toward stable symbiotic equilibria [337].